FIND THE BASE AND HEIGHT OF A TRIANGLE
area=60sq units base= 3x+6 height= x+5
area=1/2 bh help???

Answer by stanbon(75887)

FIND THE BASE AND HEIGHT OF A TRIANGLE
area=60sq units base= 3x+6 height= x+5
---------
Area = (1/2)bh
60 = (1/2)(3x+6)(x+5)
(3x+6)(x+5) = 120
3x^2 + 15x + 6x + 30 = 120
3x^2 + 21x - 90 = 0
x^2 + 7x - 30 = 0
---
x = [-7 +- sqrt(49 - 4*-30)]/2
---
x = [-7 +- sqrt(169)]/2
Positive Solution:
x = (-7 + 13]/2
x = 3
---
base = 3x+6 = 3*3+6 = 15 units
height = x+5 = 3+5 = 8 units
=================================
Cheers,
Stan H.